X=2y-5 and Y=3x-5 then X+Y=?
Answer by mathsnight(4)

Substitute the given values simply:
x=2(3x-5)-5
=> x=3; for y,
y=3x-5=3(3)-5=4.
Then x+y=3+4=7.